OpenStudy (anonymous):

f(x)=cotx For what values of x is f(x) > -√3 on the interval (0,π)?

OpenStudy (campbell_st):

well cot = 1/tan and tan is negative in the 2nd quadrant... so cot will be negative as well so \[\cot(x) = -\sqrt{3} ... then 1/\cot = \tan... so... \tan(x) = -\frac{1}{\sqrt3}\] find the value of \[x = \tan^{-1}(\frac{1}{\sqrt{3}})\] and 2nd quadrant angle will be \[\pi - x \]

OpenStudy (anonymous):

Thank you :)